it is a 42 km drive go oakdale to ridgemont and takes 28 min.
Because of traffic and construction working the trip is now 14 minutes longer.
find the reduced speed limit in km/h

Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

r=D/t
r=42/.07 hrs
r=60 km/h

Answer by richwmiller(17219)

28+14=42 minutes
42/60=7*6/10*6=7/10=.7
r*t=d
r*.7=42
7r=420
r=60
Amazingly the other tutor came up with the right answer after making two mistakes.
He used .07 which is wrong but also divided 42/.07 and got 60 which is also wrong.
42/.07=4200/7=600 But the answer is 60.